The mass of the body is M = 5 kg, while the tensile force F = 50 N forms the angle a = 30 ° with the horizontal surface with coefficient of friction = 0.15 Determine the acceleration, the normal reaction force and the compressive force on the plane
1
Expert's answer
2020-12-15T11:47:39-0500


1. Let's write down the second Newton's law in a projection to the y-axis, taking into account that the body does not move in vertical direction (acceleration "a_y = 0" ):


"N+F\sin30\degree - mg = 0"


where "F = 50N" is the tensile force, "N" is the normal reaction force, "m = 5kg" is the mass of the body, and "g = 9.8m\/s^2" is the gravitational acceleration. Expressing "N", obtain:


"N = mg - F\sin30\degree \\
N = 5\cdot 9.8- 50\cdot \sin 30\degree = 24N"



2. According to the third Newton's law, the compressive force on the plane is equal to the normal force and equal to "24N".


3. Let's write down the second Newton's law in a projection to the x-axis:


"F\cos30\degree-F_{fr} = ma"

where "F_{fr}" is the frictional force, and "a" is the acceleration of the body. By defitnion:


"F_{fr} = \mu N = 0.15\cdot 24N = 3.6N"

where "\mu =0.15" is the coefficient of friction. Thus, obtain for acceleration:


"a = \dfrac{F\cos30\degree-F_{fr}}{m} = \dfrac{50\cdot \cos30\degree-3.6}{5} \approx 7.9m/s^2"

Answer. 7.9m/s^2, 24N, 24N.